OpenStudy (anonymous):

how do you find the derivative of |x+1|-|x-3|?

OpenStudy (anonymous):

it is a piecewise function, work in cases

OpenStudy (anonymous):

for example if \(x>3\) then \(|x+1|=x+1\) and \(|x-3|=x-3\) so if \(x>3\) you have \[|x+1|-|x-3|=x+1-(x-3)=-2\] a constant the derivative of a constant is zero

OpenStudy (anonymous):

@satellite73 thats what I thought, but in for my question, it is asking for f'(2.7) of the piecewise function, and the answer given is 2 o-o

OpenStudy (anonymous):

that is because if \(-1<x<3\) you have \(|x+1|=x+1\) and \(|x-3|=3-x\)

OpenStudy (anonymous):

subtract and get \[2x-2\] and the derivative is 2

OpenStudy (anonymous):

that is the "work in cases" part

OpenStudy (anonymous):

@satellite73 why is |x−3|=3−x?

OpenStudy (anonymous):

and if a derivative of a constant is 0, why is f'(2.7) not 0 then?

OpenStudy (anonymous):

because \(2x-2\) is not a constant

OpenStudy (anonymous):

@satellite73 I am still confused about |x−3|=3−x? :s

OpenStudy (anonymous):

what is your definition of \(|x-3|\)?

OpenStudy (anonymous):

absolute value of x-3; graph is moved 3 units to the right and range is y>0

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[|x-3| = \left\{\begin{array}{rcc} -(x-3) & \text{if} & x < 3 \\ x-3& \text{if} & x \geq 3 \end{array} \right.\]

OpenStudy (anonymous):

ok so |x−3|=3−x because x<3? i think I'll just go review my absolute values unit. thanks!